OpenStudy (anonymous):

mr hough's has 30 students 2/5 play an instrument and 5/6 of them play sports. How many play sports and how many play instruments.

OpenStudy (hlares):

To figure out how many of the 30 students play sports and how many students play instruments you will need to find the product of 30 and the fraction you are using. For how many students play sports, you will first multiply 30 by the numerator, then divide that product by the denominator for your answer: 30*(5/6)=150/6=25 Thus, 25 of the 30 students play sports. For how many students play instruments, you go through the same process with the appropriate fraction: 30*(2/5)=60/5=12 Thus, 12 of the 30 students play instruments.